OpenStudy (anonymous):

find the first term and common difference of the arithmetic described. 7th term is 7, 22nd term is 52 what is the first term?

OpenStudy (cwrw238):

a + 6d = 7 where a = first term and d=common difference a + 21d = 52 (22nd term) solve this system of equations and you have it.

OpenStudy (cwrw238):

i suggest you subtract the first equation from the second - this will eliminate a

OpenStudy (anonymous):

not sure how i plug this in

OpenStudy (cwrw238):

a + 21d = 52 a + 6d = 7 -------- subtract: 15d = 45 so now you can find what d is right?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

?

OpenStudy (cwrw238):

yes d = common difference = 3 now plug d=3 into the first equation so as to find the value of the first term a.

OpenStudy (anonymous):

21*3=63

OpenStudy (cwrw238):

? a + 6d = 7 so a +6(3) = 7 a = 7 -18 a = ?

OpenStudy (anonymous):

-11

OpenStudy (cwrw238):

yup

OpenStudy (anonymous):

now i need to find an-1+?

OpenStudy (anonymous):

i got that i just need to know an?

OpenStudy (cwrw238):

? we have now found the common difference = 3 and first term which is -11

OpenStudy (anonymous):

so what is an?

OpenStudy (cwrw238):

an - that the nth term right?

OpenStudy (anonymous):

ya

OpenStudy (cwrw238):

if so it is a +(n-1)d = -11 + 3(n -1)

OpenStudy (cwrw238):

= 3n - 14

OpenStudy (anonymous):

so thats the formula than?

OpenStudy (cwrw238):

yep - that s the formula for the nth term (an) of this series